9. A 50 ml solution of unknown base contains 0.0012 mols of base (2.94 g/L) in solution.
Calculate the molar mass of the solution, and use that to calculate the concentration and
pH of the solution.
1
Expert's answer
2020-05-08T14:09:51-0400

Solution.

"V=50mL=0.05L;"

"\nu_x=0.0012mol;"

"\rho_x=2.94g/L;"

"\nu_x=\dfrac{V_x}{V_m};\implies V_x=\nu_xV_m;" "V_m" - volume of one mole;

"V_x=0.0012mol\sdot22.4L/mol=0.02688L;"

"\rho_x=\dfrac{m_x}{V_x};\implies m_x=\rho_xV_x;"

"m_x=2.94g/L\sdot0.02688L=0.079g;"

"\nu_x=\dfrac{m_x}{M_x} \implies M_x=\dfrac{m_x}{\nu_x}" ;


"M_x=\frac{0.079g}{0.0012mol}=65.856g/mol;"


"C_x=\dfrac{\nu_x}{V};" "C_x=\dfrac{0.0012mol}{0.05L}=0.024mol/L;"

"pH=-lgC_x; pH=-lg0.024=lg41.67=1.62;"

Answer:"M_x=65.856g\/mol;"

"C_x=0.024mol\/L;"

"pH=1.62;"